What is the ratio of the root-mean-squared velocity (rms) of deuterium vs. hydrogen at a given temperature? The molar mass of deuterium and hydrogen are 4.0 g/mol and 2.0 g/mol, respectively.
Answer #1

Crms = 3RT/M

C(D2)/C(H2) = M(H2)/M(D2)

molar mass of deuterium M(D2)       = 4g/mole

molar mass of hydrogen M(H2)       = 2g/mole

C(D2)/C(H2) = M(H2)/M(D2)

                   = 2/4

                   = 1/2

C(D2) : C(H2)         = 1 :1.414>>>>>answer